# An insight on some properties of high order nonstandard linear multistep methods

### v1

@@ 5 Applications and numerical simulations
In this section, we apply the previously observed nonstandard multistep method (10) to different initial value problems.
@@
    Apply the Forward Euler method to the given equation with $\Delta t$ and calculate such a bound $B_{FE}$, for which the observed property is preserved if $\Delta t\leq B_{FE}$ holds.
@@
    For a method with $s$-many steps, the values $u^{1},\;u^{2},\;\dots,\;u^{s-1}$ should be calculated. For this, a nonstandard Runge-Kutta method can be used with an appropriate function $\varphi(x)$ (see example in Section 5.2 for details).
@@ 5.1 One-dimensional logistic growth equation
Consider the one-dimensional ($n=1$) logistic growth equation
$\begin{cases}y^{\prime}(t)&=y(t)\left(2-y(t)\right),\\ y(0)&=\tilde{y}.\end{cases}$ (16)
The solution of the above equation is $y(t)=\dfrac{2\tilde{y}}{\tilde{y}+(2-\tilde{y})e^{-2x}}$. It is also apparent that the two equilibria of the equation are $y_{1}^{*}=0$ and $y_{2}^{*}=2$, where $y_{1}^{*}$ is unstable and $y_{2}^{*}$ is asymptotically stable. The solution of the equation is also monotone.
To apply the nonstandard multistep method, we should first observe the behavior of the Forward Euler method.
@@
The Forward Euler method applied to equation (16) preserves the monotonicity and the boundedness of the solution if $\Delta t\leq\min\left\{\dfrac{1}{2},\dfrac{1}{\tilde{y}}\right\}$ if $\tilde{y}\geq 0$, and if $\tilde{y}<0$, then the method preserves the properties unconditionally.
@@
$u^{n+1}=u^{n}+\Delta tu^{n}(2-u^{n})=u^{n}(1+2\Delta t-\Delta t\;u^{n}).$
@@
  • If either $\tilde{y}=0$ or $\tilde{y}=2$, the numerical solution remains constant.
@@
  • Let us assume that $\tilde{y}=u^{0}<0$. In this case, the solution should remain negative, and it should be monotone decreasing. If $u^{n}<0$, then $u^{n+1}$ is also negative if $01+2\Delta t-\Delta t\;u^{n}>$ 0, but this is true. For the decreasing property, we need that $11+2\Delta t-\Delta t\;u^{n}>$ 1, but this holds too. Consequently, in this case, the method preserves the boundedness and the monotonicity of the solution unconditionally.
@@
  • Assume that $20<\tilde{y}<$ 2. For the increasing property, it is enough to show that $11+2\Delta t-\Delta t\;u^{n}>$ 1, but this is true since $u^{n}>0$. For the boundedness from above, consider the function $g(x)=x(1+2\Delta t-\Delta t\;x)$. This attains its maximum at $\dfrac{1+2\Delta t}{2\Delta t}$, and its value at this point is $\dfrac{(1+2\Delta t)^{2}}{4\Delta t}$ which is bigger than $2$ for every positive value of $\Delta t$. Because of this, let us choose $\Delta t$ in a way that the maximum point is bigger than $2$: $\Delta t<\frac{1}{2}$ is sufficient. Therefore, the function $g$ is monotone increasing on $[0,2]$, with its maximum point at $x=2$ with value $g(2)=2$. Therefore, $u^{n+1}=g(u^{n})<2$ if $\Delta t<\dfrac{1}{2}$. The boundedness from below is trivial because of the monotone increasing property.
@@
  • Consider the case $\tilde{y}>2$. For the decreasing property we need $11+2\Delta t-\Delta t\;u^{n}<$ 1 but this is true since $u^{n}>2$. For the boundedness from below, consider the function $g(x)=x(1+2\Delta t-\Delta t\;x)$. We should prove that if $\tilde{y}>x>2$, then $g(x)>2$. In the previous case we have shown that if $\Delta t<\frac{1}{2}$, then the maximum point is bigger than $2$, so since $g(2)=2$, for a sufficiently small $\varepsilon$, $g(x)>2$ if $x\in(2,2+\varepsilon)$. Our goal is to choose $\Delta t$ in a way that $2+\varepsilon>\tilde{y}$.
@@
    If we solve $g(x)=2$, the solution (apart from the previously mentioned $x_{0}=2$) is $x_{1}=\frac{1}{\Delta t}$. Consequently, if $\Delta t\leq\frac{1}{\tilde{y}}$, then the scheme preserves the lower bound of the method. ∎
@@
In conclusion, a proper choice for parameter $\mathcal{B}$ is $\mathcal{B}:=\mathcal{C}\min\left\{\dfrac{1}{2},\dfrac{1}{\tilde{y}}\right\}$, and function $\varphi$ can be constructed in accordance with the ideas mentioned in Section 3.3. Therefore, in line with the results of Theorems 3 and 4, we expect the numerical solution produced by the nonstandard multistep method to be bounded and attain property (11) for all $\Delta t\geq 0$ step sizes. In this case, Theorem 5 is not applicable since the equation is one-dimensional (and the preservation of the constant values is trivial).
@@
In the following, we run several numerical tests observing the performance of the aforementioned numerical methods. For this, we use the strongly stable forms of multistep methods [52] - the coefficients of the methods applied are listed in the Appendix. Moreover, since the solution is known, we used the exact values for those initial values of the numerical method that are needed (for example, for an $s$-step method, one needs the values of $u^{0},\;u_{1},\;\dots,\;u^{s-1}$). Alternatively, we could have used some further, high-order (possibly nonstandard) numerical method to approximate these values - such an approach is taken in the second example in Section 5.2.

### v2

@@ 5 Applications and numerical simulations
In this section, we apply the previously analyzed nonstandard multistep method (10) to different initial value problems.
@@
    Apply the Forward Euler method to the given equation with $\Delta t$ and calculate such a bound $B_{FE}$, for which the considered property is preserved if $\Delta t\leq B_{FE}$ holds.
@@
    For a method with $s$-many steps, the values $u^{1},\;u^{2},\;\dots,\;u^{s-1}$ should be calculated. For this, a nonstandard Runge-Kutta method can be used with an appropriate function $\varphi(x)$ (see Remark 6).
@@ 5.1 One-dimensional logistic growth equation
Consider the one-dimensional ($n=1$) logistic growth equation with constant $c\in\mathbb{R}^{+}$
[…]
The solution of the above equation is $y(t)={\color[rgb]{0,0,0}\dfrac{ce^{ct}\tilde{y}}{\tilde{y}(e^{ct}-1)+c}}$. It is also apparent that the two equilibria of the equation are $y_{1}^{*}=0$ and $y_{2}^{*}=c$, where $y_{1}^{*}$ is unstable and $y_{2}^{*}$ is asymptotically stable. The solution of the equation is also monotone.
To apply the nonstandard multistep method, we should first analyze the behavior of the Forward Euler method.
@@
The Forward Euler method applied to equation (15) preserves the monotonicity and the boundedness of the solution if $\Delta t\leq\min\left\{{\color[rgb]{0,0,0}\dfrac{1}{c}},\dfrac{1}{\tilde{y}}\right\}$ and $\tilde{y}\geq 0$. If $\tilde{y}<0$, then the method preserves the properties unconditionally.
@@
$\color[rgb]{0,0,0}v^{n+1}=v^{n}+\Delta tv^{n}(c-v^{n})=v^{n}(1+c\Delta t-\Delta t\;v^{n}).$}
@@
  • If either $\tilde{y}=0$ or $\tilde{y}={\color[rgb]{0,0,0}c}$, the numerical solution remains constant.
@@
  • Let us assume that $\tilde{y}={\color[rgb]{0,0,0}v^{0}}<0$. In this case, the solution should remain negative, and it should be monotone decreasing. If ${\color[rgb]{0,0,0}v^{n}}<0$, then $\color[rgb]{0,0,0}v^{n+1}$} is also negative if $1+{\color[rgb]{0,0,0}c}\Delta t-\Delta t\;{\color[rgb]{0,0,0}v^{n}}>0$, but this is true. For the decreasing property, we need that $1+{\color[rgb]{0,0,0}c}\Delta t-\Delta t\;{\color[rgb]{0,0,0}v^{n}}>1$, but this holds too. Consequently, in this case, the method preserves the boundedness and the monotonicity of the solution unconditionally.
@@
  • Assume that $0<\tilde{y}<{\color[rgb]{0,0,0}c}$. For the increasing property, it is enough to show that $1+{\color[rgb]{0,0,0}c}\Delta t-\Delta t\;{\color[rgb]{0,0,0}v^{n}}>1$, but this is true since $\color[rgb]{0,0,0}v^{n}<c$}. For the boundedness from above, consider the function $g(x)=x(1+{\color[rgb]{0,0,0}c}\Delta t-\Delta t\;x)$. This attains its maximum at $\dfrac{1+{\color[rgb]{0,0,0}c}\Delta t}{2\Delta t}$, and its value at this point is $\dfrac{(1+{\color[rgb]{0,0,0}c}\Delta t)^{2}}{4\Delta t}$ which is bigger than $\color[rgb]{0,0,0}c$} for every positive value of $\Delta t$. Because of this, let us choose $\Delta t$ in a way that the maximum point is bigger than $\color[rgb]{0,0,0}c$}: $\Delta t<\frac{1}{{\color[rgb]{0,0,0}c}}$ is sufficient. Therefore, the function $g$ is monotone increasing on $0,{\color[rgb]{0,0,0}c}]$, with its maximum point at $x={\color[rgb]{0,0,0}c}$ with value $g({\color[rgb]{0,0,0}c})={\color[rgb]{0,0,0}c}$. Therefore, ${\color[rgb]{0,0,0}v^{n+1}=g(v^{n})}<{\color[rgb]{0,0,0}c}$ if $\Delta t<\dfrac{1}{{\color[rgb]{0,0,0}c}}$. The boundedness from below is trivial because of the monotone increasing property.
@@
  • Consider the case $\tilde{y}>{\color[rgb]{0,0,0}c}$. For the decreasing property we need $1+{\color[rgb]{0,0,0}c}\Delta t-\Delta t\;{\color[rgb]{0,0,0}v^{n}}<1$ but this is true since ${\color[rgb]{0,0,0}v^{n}}>{\color[rgb]{0,0,0}c}$. For the boundedness from below, consider the function $g(x)=x(1+{\color[rgb]{0,0,0}c}\Delta t-\Delta t\;x)$. We should prove that if $\tilde{y}>x>{\color[rgb]{0,0,0}c}$, then $g(x)>{\color[rgb]{0,0,0}c}$. In the previous case we have shown that if $\Delta t<\frac{1}{{\color[rgb]{0,0,0}c}}$, then the maximum point is bigger than $\color[rgb]{0,0,0}c$}, so since $g({\color[rgb]{0,0,0}c})={\color[rgb]{0,0,0}c}$, for a sufficiently small $\varepsilon$, $g(x)>{\color[rgb]{0,0,0}c}$ if $x\in({\color[rgb]{0,0,0}c},{\color[rgb]{0,0,0}c}+\varepsilon)$. Our goal is to choose $\Delta t$ in a way that ${\color[rgb]{0,0,0}c}+\varepsilon>\tilde{y}$.
@@
    If we solve $g(x)={\color[rgb]{0,0,0}c}$, the solution (apart from the previously mentioned $x_{0}={\color[rgb]{0,0,0}c}$) is $x_{1}=\frac{1}{\Delta t}$. Consequently, if $\Delta t\leq\frac{1}{\tilde{y}}$, then the scheme preserves the lower bound of the method. ∎
@@
In conclusion, a proper choice for parameter $\mathcal{B}$ is $\mathcal{B}:=\mathcal{C}\min\left\{\dfrac{1}{{\color[rgb]{0,0,0}c}},\dfrac{1}{\tilde{y}}\right\}$, and function $\varphi$ can be constructed in accordance with the ideas mentioned in Section 3.3. Therefore, in line with the results of Theorems 3 and 4, we expect the numerical solution produced by the nonstandard multistep method to be bounded and attain property (11) for all $\Delta t\geq 0$ step sizes. In this case, Theorem 5 is not applicable since the equation is one-dimensional (and the preservation of the constant values is trivial).
@@
In the following, we run several numerical tests analyzing the performance of the aforementioned numerical methods. For this, we use the strongly stable forms of multistep methods [54] - the coefficients of the methods applied are listed in Appendix A. Moreover, since the solution is known, we used the exact values for those initial values of the numerical method that are needed (for example, for an $s$-step method, one needs the values of $u^{0},\;u_{1},\;\dots,\;u^{s-1}$). Alternatively, we could have used some further, high-order (possibly nonstandard) numerical method to approximate these values - such an approach is taken in the second example in Section 5.2.
Initially, we examine the choice $c=2$, and then the case of $c=500$ (resulting in a strict bound for $\varphi$) is also considered.
5.1.1 The non-stiff case of $c=2$
In this subsection we consider equation (15) with $c=2$. First, we select the nonstandard version of the fourth-order multistep method SSPMS(6,4) and check the performance of the different choices of the function $\varphi(x)$. We run the methods with timesteps $\Delta t=0.1/2^{k},\;k=0,1,\dots,9$ and initial condition $\tilde{y}=1$, and compare the values of the numerical solutions at $T=1$ to the exact solution. The values of the errors are plotted in Figure 1, while the exact values of the errors and the corresponding orders can be seen in Table 2 in Appendix B. From now on, we use the notations introduced in Section 3.3. As we can see, with functions $\varphi_{1},\varphi_{2}$ and $\varphi_{3}$, the method is only of first order, with $\varphi_{4},\varphi_{5}$ and $\varphi_{6}$ we attain second order, while the uses of $\varphi_{7}$ and $\varphi_{8}$ result in a third- and fourth order methods, respectively. It is also worth mentioning here, that from the first three methods $\varphi_{2}$ performs the best, while among the second order ones, $\varphi_{5}$ seems to be the one with the smallest errors - this is somewhat anticipated since these are the ones with the smallest error terms among the methods with the same order (see the $\varphi^{\prime\prime}(0)$ and $\varphi^{\prime\prime\prime}(0)$ values in Section 3.3).
Figure 1: The errors of the different methods with different $\varphi$ functions for different values of $\Delta t$ for $c=2$.
Next, we test the order s of the nonstandard multistep methods: for this, we choose $\varphi_{8}$ (or $\varphi_{7}$, see the next paragraph) and run the nonstandard versions of methods SSPMS(4,2), SSPMS(4,3) and SSPMS(6,4) ( denoted by NSSPMS(4,2), NSSPMS(4,3) and NSSPMS(6,4)) along with the nonstandard versions of the SSP Runge-Kutta methods SSPRK(2,2), SSPRK(3,3) and SSPRK(10,4) (denoted by NSSPRK(2,2), NSSPRK(3,3) and NSSPRK(10,4), respectively). The coefficients of these methods can be found in Appendix A. Here we run the methods with timesteps $\Delta t={\color[rgb]{0,0,0}0.05/2^{k},\;k=0,1,\dots,8}$ (with $\tilde{y}=1$ as before) and compare the numerical solutions to the exact solution at $T=1$. Note that here we have three different values of $\mathcal{C}$, meaning that parameter $\mathcal{B}$ might have different values depending on the method. In Figure 2 we plot the errors, while Table 3 in Appendix B. contains the errors along with the orders. Apart from the third order ones, the methods perform as expected. In general, the multistep and the Runge-Kutta methods have a similar order, while the latter produce smaller errors.
Figure 2: The errors of the different nonstandard methods for different values of $\Delta t$ with $c=2$.
One surprising property is the order of the convergence of the third-order method NSSPMS(4,3) : its order is close to $4$ for larger values of $\Delta t$. This might be the result of the effect of the function $\varphi_{8}$, since this function enables higher order (see the cases of NSSPMS(6,4) and NSSPRK(10,4)). This is also proven by the plot of the same method run with the choice $\varphi_{7}$: in this case, since this function can only enable an order of $3$, the order of the method is decreased by one, resulting in the expected order of $3$. This might indicate that for carefully chosen $\varphi$ functions, nonstandard multistep methods might have larger orders than their standard counterparts. Consequently, it is usually advised to use such $\varphi$ functions that enable higher orders (four or even bigger). Note that this effect is not present in the Runge-Kutta case: for NSSPRK(3,3), both of the choices of $\varphi_{7}$ and $\varphi_{8}$ result in an order of $3$, although the $\varphi_{8}$ choice results in smaller errors.
Moreover, we consider the way the nonstandard methods preserve the qualitative properties of the original differential equation. By Theorems 3 and 4, we know that the nonstandard multistep method preserves the boundedness property of the original solution (since the Forward Euler method has this property for $\Delta t\leq\mathcal{B}$, see Proposition 1), and condition (11) also holds. In Figure 3, we compare the results of the nonstandard multistep methods with their standard counterparts applied with the same step sizes, initial value $\tilde{y}=3$, and function choices $\varphi_{5}$, $\varphi_{7}$, and $\varphi_{8}$, respectively. On the same figures, we also plot the solutions produced by the nonstandard Runge-Kutta methods of the same order. As it can be seen, for $\Delta t_{1}=0.5$ the standard methods are highly unstable: they are not bounded from below, and (11) is also violated. However, the nonstandard ones behave as expected, even for a larger timestep: their solutions are bounded from below, and the property (11) also holds with $s=4$ (in the first two cases) and with $s=6$ (in the last case). For $\Delta t_{2}=0.02$, all methods behave as expected. In general, the nonstandard multistep methods perform better than their standard versions, but fail to preserve e.g. monotonicity for $\Delta t_{1}$, while the nonstandard Runge-Kutta ones achieve even this goal.
Figure 3: Methods SSPMS(4,2), NSSPMS(4,2) and NSSPRK(2,2) (upper left), SSPMS(4,3), NSSPMS(4,3) and NSSPRK(3,3) (upper right) and SSPMS(6,4), NSSPMS(6,4) and NSSPRK(10,4) (lower) with timesteps $\Delta t_{1}=0.5$ and $\Delta t_{2}=0.02$. For $\Delta t_{2}$, the schemes are very close to each other, therefore, only one of them is visible.